Sayyyy10:

Two lines meet at a point that is also the vertex of a right angle. Set up and solve an equation to find the value of 𝑥. Find the measurements of ∠CAE and ∠BAG. dfg For Blank 1: What is the value of x? For Blank 2: What is the value of ∠CAE? For Blank 3: What is the value of ∠BAG ?

Sayyyy10:

us this

1 attachment
jhonyy9:

so like a first step you see that angle Toast is a right angle ,yes ? so what mean this ,angle Toast has how many degrees ? and now you can write the angle Toast like the sum of what angles ?

jhonyy9:

and look please on this image ,so how are angles <DAB and <CAE ?

surjithayer:

\[\angle FAB+\angle DAG=90^\circ\] \[\angle FAB+3x=90^\circ\] \[\angle DAF+\angle FAB=4x\] \[15+\angle FAB=4x\] \[\angle FAB=4x-15\] 4x-15+3x=90 7x=90+15 7x=105 x=105/7=15 \[\angle ~CAE=4*15=60^\circ\] \[\angle~BAG=3x=3*15=45^\circ\]

Jake15:

^